# The geometry and arithmetic of bielliptic Picard curves

### v2

@@ 7.5. $2$-torsion in $A$ using bitangents
The action of $\langle\tau\rangle$ on the $28$ bitangents of $C_{\bar{k}}$ has four fixed points (normal-(((including the line at infinity)normal-))) and $12$ orbits of size $2$.
@@ 8.1. Eliminating $\ell$-torsion for $\ell\geq 5$
We conclude by excluding the case $\ell=7$. Let $L/{\rm\mathbb{Q}}$ be the endomorphism field, namely the smallest field extension with the property that ${\rm End}(P_{L})={\rm End}(P_{\bar{{\rm\mathbb{Q}}}})$. Since ${\rm Gal}({\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}})$ acts nontrivially on the subring ${\rm\mathbb{Z}}[\omega]\subset\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$, we have ${\rm\mathbb{Q}}(\omega)\subset L$. A result of Silverberg [43, Theorem 4.2] shows that $L$ is unramified at all places of good reduction of $P$. Since $3$ is ramified in ${\rm\mathbb{Q}}(\omega)\subset L$, it follows that $P$ has bad reduction at $3$. By [28, Proposition 4.1.3(b)] and Lemma 6.9, the reduction of $P$ at $3$ is totally additive; in other words the identity component of the special fibre of the Neron model of $P$ over ${\rm\mathbb{Z}}_{3}$ is unipotent. A result of Lorenzini [32, Corollary 3.25] then shows that $7\nmid|P({\rm\mathbb{Q}}_{3})_{\mathrm{tors}}|$, a contradiction. ∎

### v3

@@ 7.5. $2$-torsion in $A$ using bitangents
The action of $\langle\tau\rangle$ on the $28$ bitangents of $C_{\bar{k}}$ has four fixed points ((((including the line at infinity)))) and $12$ orbits of size $2$.
@@ 8.1. Eliminating $\ell$-torsion for $\ell\geq 5$
We conclude by excluding the case $\ell=7$. Let $L/{\rm\mathbb{Q}}$ be the endomorphism field, namely the smallest field extension with the property that ${\rm End}(P_{L})={\rm End}(P_{\bar{{\rm\mathbb{Q}}}})$. Since ${\rm Gal}({\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}})$ acts nontrivially on the subring ${\rm\mathbb{Z}}[\omega]\subset\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$, we have ${\rm\mathbb{Q}}(\omega)\subset L$. A result of Silverberg [44, Theorem 4.2] shows that $L$ is unramified at all places of good reduction of $P$. Since $3$ is ramified in ${\rm\mathbb{Q}}(\omega)\subset L$, it follows that $P$ has bad reduction at $3$. By [28, Proposition 4.1.3(b)] and Lemma 6.9, the reduction of $P$ at $3$ is totally additive; in other words the identity component of the special fibre of the Neron model of $P$ over ${\rm\mathbb{Z}}_{3}$ is unipotent. A result of Lorenzini [32, Corollary 3.25] then shows that $7\nmid|P({\rm\mathbb{Q}}_{3})_{\mathrm{tors}}|$, a contradiction. ∎
@@ 8.3. Eliminating points of order $4$
We do not classify all the possibilities for $J({\rm\mathbb{Q}})_{\mathrm{tors}}$, but Theorem 1.1 quickly implies the following:
Corollary 8.7.
Let $C/{\rm\mathbb{Q}}$ be a bielliptic Picard curve with Jacobian $J$. Then $J({\rm\mathbb{Q}})_{\mathrm{tors}}=J({\rm\mathbb{Q}})[12]$.
Proof.
Using the notation of §2.5, the surjective maps $J\rightarrow A$ and $J\rightarrow E$ combine to an isogeny $\theta\colon J\rightarrow A\times E$ whose kernel is isomorphic to $E[2]$. Theorem 1.1 and bigonal duality (Proposition 2.19) show that $A({\rm\mathbb{Q}})_{\mathrm{tors}}$ is $6$-torsion. Moreover elementary arguments (using reduction mod $3$ and $5$) show that $E({\rm\mathbb{Q}})_{\mathrm{tors}}$ is $6$-torsion. It follows that for every $x\in J({\rm\mathbb{Q}})_{\mathrm{tors}}$, $06\theta(x)=$ 0, so $6x\in\ker(\theta)\simeq E[2]$, which is $2$-torsion. ∎